What is the solution?


A man holds a ball of weight 114g at rest in his hand. He then throws the ball vertically upward. In this process, his hand moves up 0.6m and the ball leaves his hand with an upward velocity of 15 m/s. Find the force with which the man pushes on the ball.


Answer = 22.5 N


1
Expert's answer
2021-06-25T05:46:02-0400

The kinetic energy of the ball is


"E_K=\dfrac{mv^2}{2}"

"E_K=\dfrac{0.114\ kg(15\ m/s)^2}{2}=12.825\ J"

Then


"E_K=Fs=>F=\dfrac{E_K}{s}"

"F=\dfrac{12.825\ J}{0.6\ m}=21.375\ N"